# Trainable and Explainable Simplicial Map Neural Networks

### 11 Mar 2024 v2

@@ Abstract
Simplicial map neural networks (SMNNs) are topology-based neural networks with interesting properties such as universal approximation ability and robustness to adversarial examples under appropriate conditions. However, SMNNs present some bottlenecks for their possible application in high-dimensional datasets. First, SMNNs have precomputed fixed weight and no SMNN training process has been defined so far, so they lack generalization ability. Second, SMNNs require the construction of a convex polytope surrounding the input dataset. In this paper, we overcome these issues by proposing an SMNN training procedure based on a support subset of the given dataset and replacing the construction of the convex polytope by a method based on projections to a hypersphere. In addition, the explainability capacity of SMNNs and an effective implementation is also newly introduced in this paper.
@@ 2.4 Simplicial map neural networks
In the original definition [5], the first hidden layer of an SMNN computes the barycentric coordinates of $\operatorname{Del}(V\cup P)$.
However, we will see here that if we precompute the barycentric coordinates, we can simplify the architecture of $\mathcal{N}_{\varphi}$ as follows.
@@
satisfying that, for $t\in[\![1,\alpha]\!]$, $\xi_{t}(x)=b_{j}(x)$ if $i_{j}=t$ for some $j\in[\![0,n]\!]$.
@@ 3 The unknown boundary and the function $\varphi_{\scriptscriptstyle U}$
First, let us compute a hypersphere surrounding $V$. One of the simplest ways to do that is to translate $V$ so that its centre of mass is placed at the origin $o\in\mathbb{R}^{n}$. Then, the hypersphere
@@
Now, let us define ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\xi_{U}}}(x)\in\mathbb{R}^{m}$ for any $x\in B^{n}$ as follows. Given $x\in B^{n}$, to find the $n$-simplex $\sigma=\langle\omega^{0},\dots,\omega^{n}\rangle$ such that $x\in|\sigma|$, we have to consider two cases: $x\in|\operatorname{Del}(U)|$ and $x\not\in|\operatorname{Del}(U)|$.
@@
Now, let $b(x)=(b_{0}(x),\dots,b_{n}(x))\in\mathbb{R}^{n+1}$ be the barycentric coordinates of $x$ with respect to $\sigma$. Then, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\xi_{U}}}(x)=(\xi_{1}(x),\dots,\xi_{m}(x))$ is the point in $\mathbb{R}^{m}$ satisfying that, for $t\in[\![1,m]\!]$,
@@
Observe that ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\xi_{U}}}(x)$ always exists and is unique. An example of points $x$ and $w^{x}$ and simplex $\mu$ is shown in Figure 3 and Example 1.
Let us observe that, thanks to the new definition of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\xi_{U}}}(x)$ for $x\in B^{n}$, if we have a map $\varphi_{\scriptscriptstyle U}^{\scriptscriptstyle(0)}:U\to\mathbb{R}^{k}$ then it induces a continuous function $\varphi_{\scriptscriptstyle U}:B^{n}\to|L|$ defined for any $x\in B^{n}$ as:
$\begin{array}[]{cl}\varphi_{\scriptscriptstyle U}(x)&=\operatorname{softmax}\big{(}\sum_{\scriptscriptstyle t\in[\![1,m]\!]}\xi_{t}(x)\varphi_{\scriptscriptstyle U}^{\scriptscriptstyle(0)}(u^{t})\big{)}\\ &{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{=\operatorname{softmax}(\mathcal{M}_{U}\cdot\xi_{U}(x))}}\end{array}$
@@
with label $\lambda^{1}=0$ for $v^{i}$ with $i=1$ and $\lambda^{2}=1$ for $v^{i}$ with $i=2,3,4$. Firstly, we translate $V$ so that the center of mass of $V$ is the origin $o\in\mathbb{R}^{2}$. The translated dataset $\tilde{V}$ is
@@
Let us consider $x^{1}=\big{(}\frac{3}{4},\frac{3}{5}\big{)}$ and $x^{2}=\big{(}\frac{3}{4},\frac{5}{4}\big{)}.$ Hence, the translated input data is $\tilde{x}^{1}=\big{(}0,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\frac{-3}{20}}}\big{)}$ and $\tilde{x}^{2}=\big{(}0,\frac{1}{2}\big{)}$.
@@
  • Since the barycentric coordinates of $\tilde{x}^{1}$ with respect to $|\tilde{\sigma}^{1}|$ are $(0.5,0.3,0.2)$ then $\tilde{x}^{1}$ is in $|\tilde{\sigma}^{1}|\subset|\operatorname{Del}(U)|$ and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\xi_{U}}}(\tilde{x}^{1})=(0.3,0.2,0,0.5)$. Then
    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\varphi_{U}}}(\tilde{x}^{1})=\operatorname{softmax}(\mathcal{M}_{U}\cdot\xi_{U}(\tilde{x}^{1}))=(0.5,0.5).$
  • On the other hand, the point $\tilde{x}^{2}$ is outside $|\operatorname{Del}(U)|$. Assuming that, for example, we have fixed $R=1$, we add a new simplex $\sigma^{3}=\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{w^{x}}},\tilde{v}^{1},\tilde{v}^{2}\}$ where ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{w^{x}=}}\tilde{v}^{5}=(0,1)$ which is the projection of $\tilde{x}^{2}$ to the hypersphere of radius $R$ centered in the origin. See Figure 4. Then, the barycentric coordinates of $\tilde{x}^{2}$ with respect to $\sigma^{3}$ is $(0.33,0.33,0.33)$ and then ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\xi_{U}}}(\tilde{x}^{2})=(0,0.33,0.33,0)$, concluding that
    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\varphi_{U}}}(\tilde{x}^{2})=\operatorname{softmax}(\mathcal{M}_{U}\cdot\xi_{U}(\tilde{x}^{2}))=(0.5,0.5).$
@@
The pseudocode for computing $\varphi_{U}(x)$ is provide in Algorithm 1.
@@
     compute $b(x)$ with respect to $\sigma$ abd $\xi_{U}(x)$
@@
$\lim_{y\to x}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\xi_{U}}}(y)={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\xi_{U}}}(x).$
Proof. If $x\in|\operatorname{Del}(U)|$, then the result holds due to the continuity of the barycentric coordinates transformation. If $x\not\in|\operatorname{Del}(U)|$, since the origin $o\in|\operatorname{Del}(U)|$, then $||x||\neq 0$. Therefore, for $y$ close to $x$, $||y||\neq 0$ and $w^{y}=R\frac{y}{||y||}\in\mathbb{R}^{n}$. Besides,
@@
$\lim_{y\to x}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\xi_{U}}}(y)={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\xi_{U}}}(x)\ ,$
@@ 4 Training SMNNs
Let us now see how we add trainability to the SMNN $\mathcal{N}_{\varphi_{\scriptscriptstyle U}}$ induced by $\varphi_{\scriptscriptstyle U}$. Let $V$ be the training set and let $U$ be a support set lying in the same space than $V$. First, assuming that $U=\{u^{1},\dots,u^{m}\}$ has $m$ elements, then $\mathcal{N}_{\varphi_{\scriptscriptstyle U}}$ is a multiclass perceptron with an input layer with $m$ neurons that predicts the $h$-th label for $h\in[\![1,k]\!]$ using the formula:
$\mathcal{N}_{\varphi_{\scriptscriptstyle U}}(x)=\operatorname{softmax}\big{(}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\tilde{\mathcal{M}}}}\cdot{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\xi_{U}}}(x)\big{)}$
where ~${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\tilde{\mathcal{M}}}}=(p_{j}^{t})_{j\in[\![1,k]\!],\,t\in[\![1.m]\!]}$ is a matrix of weights and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\xi_{U}}}(x)\in\mathbb{R}^{m}$ is obtained from the barycentric coordinates of $x\in B^{n}$ as in Section 3. Let us observe that
$\operatorname{softmax}\big{(}\tilde{\mathcal{M}}\cdot{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\xi_{U}}}(x)\big{)}\in|L|.$
@@
in order to obtain new values for ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathcal{N}_{\varphi_{\scriptscriptstyle U}}}}(v)$ for $v\in V$ in a way that the error $\mathcal{L}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathcal{N}_{\varphi_{\scriptscriptstyle U}},}}\varphi,v)$ decreases. We will do it by avoiding recomputing $\operatorname{Del}(U)$ or the barycentric coordinates $(b_{0}(v),\dots,b_{n}(v))$ for each $v\in V$ during the training process.
In this way, given $v\in V$, if $v\in|\operatorname{Del}(U)|$, we compute the maximal simplex $\sigma=\langle u^{i_{0}},\dots,u^{i_{n}}\rangle\in\operatorname{Del}(U)$ such that $v\in|\sigma|$ and $i_{h}\in[\![1,m]\!]$ for $h\in[\![0,n]\!]$. If $v\not\in|\operatorname{Del}(U)|$, we compute $w\in S^{n}$ and the simplex $\sigma=\langle w,u^{i_{1}},\dots,u^{i_{n}}\rangle$ such that $v\in|\sigma|$ and $i_{h}\in[\![1,m]\!]$ for $h\in[\![1,n]\!]$. Then we compute the barycentric coordinates $b(v)$ of $v$ with respect to $\sigma$ and the point ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\xi_{U}}}(x)=(\xi_{1}(x),\dots,\xi_{m}(x))\in\mathbb{R}^{m}$ as in Section 3.
@@
$p_{j}^{t}:=p_{j}^{t}-\eta\frac{\partial\mathcal{L}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathcal{N}_{\varphi_{U}},}}\varphi,v)}{\partial p_{j}^{t}}=p_{j}^{t}-\eta(s_{j}-y_{j})\xi_{t}(v)$.
@@
     compute ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\xi_{U}}}(v)$ as in Section 3
@@
         $p:=p-\eta\frac{\partial\mathcal{L}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathcal{N}_{\varphi_{U}},}}\varphi,v)}{\partial p}$
@@
$\mathcal{N}_{\varphi_{U}}(x):=\operatorname{softmax}(\tilde{\mathcal{M}}\cdot{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\xi_{U}}}(x))$

### 12 Mar 2024 v3

@@ Abstract
Simplicial map neural networks (SMNNs) are topology-based neural networks with interesting properties such as universal approximation ability and robustness to adversarial examples under appropriate conditions. However, SMNNs present some bottlenecks for their possible application in high-dimensional datasets. First, SMNNs have precomputed fixed weight and no SMNN training process has been defined so far, so they lack generalization ability. Second, SMNNs require the construction of a convex polytope surrounding the input dataset. In this paper, we overcome these issues by proposing an SMNN training procedure based on a support subset of the given dataset and replacing the construction of the convex polytope by a method based on projections to a hypersphere. In addition, the explainability capacity of SMNNs and effective implementation are also newly introduced in this paper.
@@ 2.4 Simplicial map neural networks
In the original definition [5], the first hidden layer of an SMNN computes the barycentric coordinates of $\operatorname{Del}(V\cup P)$. However, we will see here that if we precompute the barycentric coordinates, we can simplify the architecture of $\mathcal{N}_{\varphi}$ as follows.
@@
satisfying that, for $t\in[\![1,\alpha]\!]$, $\xi_{t}(x)=b_{j}(x)$ if $i_{j}=t$ for some $j\in[\![0,n]\!]$. Let us remark that $\xi(x)$ should be a column vector, but we will use row notation, for simplicity.
@@ 3 The unknown boundary and the function $\varphi_{\scriptscriptstyle U}$
First, let us compute a hypersphere surrounding $V$. One of the simplest ways to do that is to translate $V$ so that its center of mass is placed at the origin $o\in\mathbb{R}^{n}$. Then, the hypersphere
@@
Now, let us define $\xi_{U}(x)\in\mathbb{R}^{m}$ for any $x\in B^{n}$ as follows. Given $x\in B^{n}$, to find the $n$-simplex $\sigma=\langle\omega^{0},\dots,\omega^{n}\rangle$ such that $x\in|\sigma|$, we have to consider two cases: $x\in|\operatorname{Del}(U)|$ and $x\not\in|\operatorname{Del}(U)|$.
@@
Now, let $b(x)=(b_{0}(x),\dots,b_{n}(x))\in\mathbb{R}^{n+1}$ be the barycentric coordinates of $x$ with respect to $\sigma$. Then, $\xi_{U}(x)=(\xi_{1}(x),\dots,\xi_{m}(x))$ is the point in $\mathbb{R}^{m}$ satisfying that, for $t\in[\![1,m]\!]$,
@@
Observe that $\xi_{U}(x)$ always exists and is unique. An example of points $x$ and $w^{x}$ and simplex $\mu$ is shown in Figure 3 and Example 1.
Let us observe that, thanks to the new definition of $\xi_{U}(x)$ for $x\in B^{n}$, if we have a map $\varphi_{\scriptscriptstyle U}^{\scriptscriptstyle(0)}:U\to\mathbb{R}^{k}$ then it induces a continuous function $\varphi_{\scriptscriptstyle U}:B^{n}\to|L|$ defined for any $x\in B^{n}$ as:
$\begin{array}[]{cl}\varphi_{\scriptscriptstyle U}(x)&=\operatorname{softmax}\big{(}\sum_{\scriptscriptstyle t\in[\![1,m]\!]}\xi_{t}(x)\varphi_{\scriptscriptstyle U}^{\scriptscriptstyle(0)}(u^{t})\big{)}\\ &=\operatorname{softmax}(\mathcal{M}_{U}\cdot\xi_{U}(x))\end{array}$
@@
with label $\lambda^{1}=0$ for $v^{i}$ with $i=1,2$ and $\lambda^{2}=1$ for $v^{i}$ with $i=3,4$. Firstly, we translate $V$ so that the center of mass of $V$ is the origin $o\in\mathbb{R}^{2}$. The translated dataset $\tilde{V}$ is
@@
Let us consider $x^{1}=\big{(}\frac{3}{4},\frac{3}{5}\big{)}$ and $x^{2}=\big{(}\frac{3}{4},\frac{5}{4}\big{)}.$ Hence, the translated input data is $\tilde{x}^{1}=\big{(}0,\frac{-3}{20}\big{)}$ and $\tilde{x}^{2}=\big{(}0,\frac{1}{2}\big{)}$.
@@
  • Since the barycentric coordinates of $\tilde{x}^{1}$ with respect to $|\tilde{\sigma}^{1}|$ are $(0.5,0.3,0.2)$ then $\tilde{x}^{1}$ is in $|\tilde{\sigma}^{1}|\subset|\operatorname{Del}(U)|$ and $\xi_{U}(\tilde{x}^{1})=(0.3,0.2,0,0.5)$. Then
    $\varphi_{U}(\tilde{x}^{1})=\operatorname{softmax}(\mathcal{M}_{U}\cdot\xi_{U}(\tilde{x}^{1}))=(0.5,0.5).$
  • On the other hand, the point $\tilde{x}^{2}$ is outside $|\operatorname{Del}(U)|$. Assuming that, for example, we have fixed $R=1$, we add a new simplex $\sigma^{3}=\{w^{x},\tilde{v}^{1},\tilde{v}^{2}\}$ where $w^{x}=\tilde{v}^{5}=(0,1)$ which is the projection of $\tilde{x}^{2}$ to the hypersphere of radius $R$ centered in the origin. See Figure 4. Then, the barycentric coordinates of $\tilde{x}^{2}$ with respect to $\sigma^{3}$ is $(0.33,0.33,0.33)$ and then $\xi_{U}(\tilde{x}^{2})=(0,0.33,0.33,0)$, concluding that
    $\varphi_{U}(\tilde{x}^{2})=\operatorname{softmax}(\mathcal{M}_{U}\cdot\xi_{U}(\tilde{x}^{2}))=(0.5,0.5).$
@@
The pseudocode for computing $\varphi_{U}(x)$ is provided in Algorithm 1.
@@
     compute $b(x)$ with respect to $\sigma$ and $\xi_{U}(x)$
@@
$\lim_{y\to x}\xi_{U}(y)=\xi_{U}(x).$
Proof. If $x\in|\operatorname{Del}(U)|$ then the result holds due to the continuity of the barycentric coordinates transformation. If $x\not\in|\operatorname{Del}(U)|$, since the origin $o\in|\operatorname{Del}(U)|$, then $||x||\neq 0$. Therefore, for $y$ close to $x$, $||y||\neq 0$ and $w^{y}=R\frac{y}{||y||}\in\mathbb{R}^{n}$. Besides,
@@
$\lim_{y\to x}\xi_{U}(y)=\xi_{U}(x)\ ,$
@@ 4 Training SMNNs
Let us now see how we add trainability to the SMNN $\mathcal{N}_{\varphi_{\scriptscriptstyle U}}$ induced by $\varphi_{\scriptscriptstyle U}$. Let $V$ be the training set and let $U$ be a support set lying in the same space as $V$. First, assuming that $U=\{u^{1},\dots,u^{m}\}$ has $m$ elements, then $\mathcal{N}_{\varphi_{\scriptscriptstyle U}}$ is a multiclass perceptron with an input layer with $m$ neurons that predicts the $h$-th label for $h\in[\![1,k]\!]$ using the formula:
$\mathcal{N}_{\varphi_{\scriptscriptstyle U}}(x)=\operatorname{softmax}\big{(}\tilde{\mathcal{M}}\cdot\xi_{U}(x)\big{)}$
where ~$\tilde{\mathcal{M}}=(p_{j}^{t})_{j\in[\![1,k]\!],\,t\in[\![1.m]\!]}$ is a matrix of weights and $\xi_{U}(x)\in\mathbb{R}^{m}$ is obtained from the barycentric coordinates of $x\in B^{n}$ as in Section 3. Let us observe that
$\operatorname{softmax}\big{(}\tilde{\mathcal{M}}\cdot\xi_{U}(x)\big{)}\in|L|.$
@@
in order to obtain new values for $\mathcal{N}_{\varphi_{\scriptscriptstyle U}}(v)$ for $v\in V$ in a way that the error $\mathcal{L}(\mathcal{N}_{\varphi_{\scriptscriptstyle U}},\varphi,v)$ decreases. We will do it by avoiding recomputing $\operatorname{Del}(U)$ or the barycentric coordinates $(b_{0}(v),\dots,b_{n}(v))$ for each $v\in V$ during the training process.
In this way, given $v\in V$, if $v\in|\operatorname{Del}(U)|$, we compute the maximal simplex $\sigma=\langle u^{i_{0}},\dots,u^{i_{n}}\rangle\in\operatorname{Del}(U)$ such that $v\in|\sigma|$ and $i_{h}\in[\![1,m]\!]$ for $h\in[\![0,n]\!]$. If $v\not\in|\operatorname{Del}(U)|$, we compute $w\in S^{n}$ and the simplex $\sigma=\langle w,u^{i_{1}},\dots,u^{i_{n}}\rangle$ such that $v\in|\sigma|$ and $i_{h}\in[\![1,m]\!]$ for $h\in[\![1,n]\!]$. Then we compute the barycentric coordinates $b(v)$ of $v$ with respect to $\sigma$ and the point $\xi_{U}(x)=(\xi_{1}(x),\dots,\xi_{m}(x))\in\mathbb{R}^{m}$ as in Section 3.
@@
$p_{j}^{t}:=p_{j}^{t}-\eta\frac{\partial\mathcal{L}(\mathcal{N}_{\varphi_{U},}\varphi,v)}{\partial p_{j}^{t}}=p_{j}^{t}-\eta(s_{j}-y_{j})\xi_{t}(v)$.
@@
     compute $\xi_{U}(v)$ as in Section 3
@@
         $p:=p-\eta\frac{\partial\mathcal{L}(\mathcal{N}_{\varphi_{U}},\varphi,v)}{\partial p}$
@@
$\mathcal{N}_{\varphi_{U}}(x):=\operatorname{softmax}(\tilde{\mathcal{M}}\cdot\xi_{U}(x))$